My son is having a problem with one specific algebra problem. I have not done this in many years. Could you help? The problem is : The trees are fruit trees and today we collected 26 pieces of fruit. There were six more apples than pears. How many of each fruit did we collect?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

The trees are fruit trees and today we collected 26 pieces of fruit. There were six more apples than pears. How many of each fruit did we collect?
-------
A + P = 26
A - P = 6
----------------- Add

Answer by ikleyn(52794)

.
Children in a kindergarten were not able to solve this problem.


Then Santa Claus approached them. He took 6 apples from the basket and put them aside.



Then he asked the children: can you solve it now ?



The smartest girl in the group answered:  


    Sure ! Now there is equal number of apples and pears in the basket, and the total is  (16-6) + 10 = 10 + 10 = 20.

           So, there are exactly 10 apples and 10 pears in the basket now.



Great! - Santa Claus said. - Now add that 6 apples !



Then the smartest boy in the group said:

    - I know the answer !  16 apples and 10 pears !



Good job ! - Santa Claus said.


He gave 1 cookie to each child from his bag and continued his way . . . 


-----------------

If your son is pre-Algebra, then consider THIS solution:

Let "x" be the number of pears.

Then the number of apples is  (x+6), according to the condition.


The total number of fruits is


    x + (x+6) = 26.


It is your basic equation,

and as soon as you got it, you just completed the setup.


Now simplify and solve it:


    2x + 6 = 26

    2x = 26 - 6

    2x = 20

    x = 20/2 = 10  pears.                    ANSWER

Then the number of apples is  10+6 = 16.     ANSWER

Solved.